OpenStudy (howard-wolowitz):

1. In a statistics course, a linear regression equation was computed to predict the final exam score from the score on the first test. The equation was yˆ=10+.9x where yˆ is the final exam score and x is the score on the first test. If Shoshana scored 95 on the first test, what is the predicted value of her score on the final exam? (a) 95 (b) 85.5 (c) 90 (d) 95.5 I have to show all work-so i can get credit

OpenStudy (howard-wolowitz):

@Michele_Laino

OpenStudy (michele_laino):

I think that we have to replace \(x=95\) into your formula above

OpenStudy (howard-wolowitz):

to get the answer

OpenStudy (michele_laino):

yes!

OpenStudy (howard-wolowitz):

what about y

OpenStudy (michele_laino):

here is the computation: \(\hat y=10+ 0.9 \cdot 95=...?\)

OpenStudy (howard-wolowitz):

95.5

OpenStudy (michele_laino):

that's right!

OpenStudy (howard-wolowitz):

sweet

OpenStudy (michele_laino):

:)

OpenStudy (howard-wolowitz):

2. Refer to problem #1. On the final exam Shoshana scores a 98. What is the value of her residual? (a) 98 (b) 2.5 (c) -2.5 (d) 0 im sorry but this problem is about the first one

OpenStudy (michele_laino):

The requested resiadual value is: \(98-95.5=...?\)

OpenStudy (michele_laino):

oops..residual*

OpenStudy (howard-wolowitz):

2.5 = B

OpenStudy (michele_laino):

by definition, the residual value is given by the difference between the observed value and the corresponding predicted value

OpenStudy (michele_laino):

that's right!